OpenStudy (anonymous):

Part 1: Explain, in complete sentences, how you would use the elimination method to solve the following system of equations. (4 points) Part 2: Provide the solution to the system. (2 points) 8x + 5y = 3 7x – 3y = 10

OpenStudy (anonymous):

Do you not understand the elimination method?

OpenStudy (anonymous):

no i hate this

OpenStudy (anonymous):

part of algebra

OpenStudy (anonymous):

What part of it do you not understand?

OpenStudy (anonymous):

this whole section

OpenStudy (anonymous):

Equations in a system can be added/subtracted from one another. Also, any individual equation can be multiplied/divided by any number without changing the graph of the equation. The elimination method is where you multiply/divide one or both of the equations with the goal of eliminating one of the variables (x or y) when you add or subtract the two equations together. Once one of the variables is eliminated, you can simply solve for the remaining variable's value.

OpenStudy (anonymous):

So looking at this one, what would you get if you simply subtracted the 2nd equation from the first equation?

OpenStudy (anonymous):

multiply first equation by 3 and the second by 5 this will allow us to eliminate y: 8x + 5y = 3 - * by 3 7x – 3y = 10 - * by 5 so we get 24x + 15y = 9 35x - 15y = 50 add these 2 equations: 59x = 59 x = 1 now plug x=1 into the first equation and solve for y